# Large Language Models in Qualitative Research: Can We Do the Data Justice?

### v1

@@ 4. Findings
In the findings section, we dive into how the researchers we interviewed that adopted LLMs into their qualitative research processes, we identify the main risks and concerns they raised, and conclude by outlining the fundamental philosophical tensions that arise in considering how qualitative work may be combined with LLMs.
@@ 4.1. Current and Potential Uses of Qualitative Researchers Adopting LLMs
We sought to understand how our participants had experimented, or not, with using LLMs as part of their qualitative research process. Of the researchers we interviewed, fourteen reported having actively explored the use of LLMs in data collection or analysis. The most common use cases mentioned, highlighted in Figure 1, were using LLMs to help generate recruitment materials, attempts to use LLMs to speed up qualitative coding, and using LLM to inspire or provide feedback on ideas. In this section, we foreground perceived potential uses of these methods, before we platform the perceptions and risks of their use in the next.
Unsurprisingly, ChatGPT was the most frequently mentioned LLM tool by participants. When we opened the conversation by asking participants about their experiences with “generative AI tools or large language models,” most would immediately talk about ChatGPT. Some researchers specified that there is a distinction; for example, P11 opened by saying “I use LLMs, specifically ChatGPT.” Others spoke more hesitantly, such as P13: “I’ve started using ChatGPT, I’m trying to be more familiar with it because it is the future, and you don’t want to be a dinosaur.”
Outside of ChatGPT, researchers noted that they used a suite of available tools for qualitative research, such as NVivo, Atlas.ti, or Dedoose, and AI sometimes underlies these tools. Within these tools, participants were sometimes unsure whether they were leveraging LLM-based functionality. For example, participants (P4, P8, P9, P13) wondered: “does Grammarly count?” Others mentioned explicitly that they engaged in AI features offered by these platforms, such as P7, who explained that through the tool Atlas.ti “people are getting suggestions […] in terms of coding certain paragraphs or sentences as in our text or document.” Participants did not necessarily feel agency over the intelligence embedded in these tools, as P1 explained: “Atlas sends me a thematic analysis that essentially aggregates stuff, I don’t really like it,” and P4 summarized AI features in these tools as “not by choice.” Overall, participants were used to relying on technology platforms to help with research, though not all appreciated the increased inclusion of smart features.
In contrast, some researchers mentioned centering a purposeful lack of advanced technology in their qualitative process. While all participants used their computers for analysis, they also talked about paring back their approach to remain close to the data. P6 explained that while they were “not quite as old school as the people who would cut out their coded texts and move it around, I basically do that in Word documents. […] I don’t even use a qualitative management software.” A similar sentiment was echoed by P2, who recounted “I didn’t like the restrictions and things the platform was putting on me, and I got really frustrated; I’m just familiar with Excel.” These attitudes highlight how, for some researchers, any software interferes with their process.
@@ 4.1.1. LLMs during Data Collection
Seven participants reported using LLMs to some degree to help with their recruitment or interview materials. Often, this took the form of researchers using a service like ChatGPT to edit or tweak their writing. Even those who were reluctant to use LLMs for research conceded that they could be helpful for writing, such as P10 who said “the only real reason that I use ChatGPT is to have it reword or rephrase,” and P4 similarly explained that “it will give me a really good sentence or switch a word.” This capability was particularly helpful for participants whose native language is not English; for example, P8 said they “feel more confident once I have passed my writing to the AI tools.” LLMs also assisted participants in tailoring recruitment materials to their subject population. One notable example was P9, a non-American researcher whose first language is not English, who told us they used LLMs to make messaging more culturally sensitive to research participants. They shared, “one thing I noticed that LLMs can do beautifully is make things American,” and therefore “an LLM’s general answer or general email to an American future participant will probably be better than my take on it.” P5 reported that LLMs helped them deploy recruitment materials in multiple languages, explaining that they “used it for the translation of interviews and consent forms and other required documents,” before having a native speaking collaborator double check the output. LLMs were also used by participants to help them create interview and recruitment materials from scratch. For example, P8 recalled “I was working with a deadline, and I was feeling stuck. I was not sure what kinds of questions to ask, but I did have a set of broad topics […], so I put those into ChatGPT and I was like, ‘Can you help me come up with this type of interview question?”’ While our participants had experimented with using these tools to help generate research materials, they did not all find it helpful: “the performance was really great,” P5 said, while P8 thought “I feel like I could have done better,” and P10 elaborated that outputs were “really broad, versus the empirical work I’ve done is very specific and granular.”
@@
In terms of promising potential uses, researchers seemed to understand why one might use LLMs to help create recruitment materials, even when they did not use them themselves. Some experienced participants explained that they had their own past templates, but P10 added “I can imagine that maybe using something like this would be helpful for beginner researchers.” P20 expressed a similar feeling: “I would consider it in the future, if I’m drafting a new consent letters or things like that, but we had all of that from previous ones.” Nonetheless, some researchers pointed to specific reasons why they may not consider LLMs always to be appropriate for these use cases, such as when materials are being produced for at-risk or non-Western populations, which will be further discussed in the following sections
@@ 4.1.2. LLMs during Data Analysis
Another avenue that participants had explored, though often less successfully, was using an LLM to make process and make sense of their qualitative data. Many researchers we talked to did not attempt such analysis, since they struggled to understand whether they could share even anonymized participant data with third-party LLMs, an issue we explore in depth in the next section. Among those interested in LLM-facilitated analysis of qualitative data, we saw a few functions being performed with LLMs: transforming or retrieving data, data exploration, generating initial codes or themes, and applying qualitative codes.
@@
Researchers wanted to use LLMs as assistants for retrieving, processing, transforming, or reformatting their data using natural language. P9 reported that they leverage LLMs to perform fuzzy searches on qualitative data: “With retrieval via an LLM, I can say something like, ‘Where in this document is an example of such and such?’ and just talk more conceptually or more abstractly, and the tool will find it.” Participants also saw LLMs as being potential useful to reformat data, for example P4 imagined that “it would be nice if there was like a large language model agent to grab all the qualitative codes [and put them] onto a Google Jamboard because I have to rely on copy pasting.”
@@
On the analysis front, participants explained that they valued being able to use LLM-aided software to initially explore participant data. In P15’s words, the ability to easily analyze qualitative data with a sensemaking model was a “game changer,” since “there’s no friction between the person trying to make sense of a lot of data and the information.” One participant said they once used ChatGPT to help them make sense of the theory behind their qualitative data, explaining that they had interviewed people about their experiences with grief. They were “getting all this lovely, well-gelled data where people were repeatedly saying the same things, everything was perfectly categorized” but were “having a hard time naming those categories in a way that was as inclusive as it could be.” In this example, the participant asked ChatGPT to come up with a list of potential themes, based not on their participant data but on prior theories, and it gave them “this list of things, and I was like, hmm, this is really neatly packaged […] it’s giving me labels for the things I’m seeing that are also very intuitive.” These themes then served as a jumping off point for deeper analysis.
@@
Experiences were more mixed for participants who directly tried to use ChatGPT or other LLMs to assign qualitative codes to data. The most common experience was perhaps that of P5, who said “I tried to use ChatGPT to get the very general theme and the results were not that great, so I decided to code it manually, but I tried it.” P9 engages in a process they refer to as “semi-coding,” where they provide ChatGPT a chunk of text they already had intuitions about, and wanted to see if the LLM agreed with their perspective. Ultimately, however, they felt that the codes were “uninteresting in a way, or at least they were first level codes.” This experience is also shared by P11, who explained that when they “asked it to assign some initial codes [on their dataset], oftentimes the response will be very high level and abstract, so it’s not trustworthy at all.” P17 disagreed, and had set up specific coding infrastructure that allowed them to identify whether an individual piece of social media data contains a rare occurrence of interest to their research. P17 was seeking to identify and code specific types of gender-based online harassment in Reddit comments and derived the most advanced annotation pipeline we saw:
@@
In their workflow, P17 derived codes based on initial data, and then constructed an LLM pipeline to annotate each post, using the LLM to take a first cut at organizing the data. P17 expressed that this process does not preclude deep qualitative work, since they still do a manual layer of qualitative analysis.
Despite overall mixed current experiences stemming from perceptions of ChatGPT, and privacy concerns with today’s systems, many participants felt curious or excited by the prospect of using LLMs to automate qualitative coding. The biggest factor that motivated curiosity around the use of LLMs for annotation was the tedium and time investment required by manual coding. P10 said “I’m currently in the process of coding, and sometimes I’m kind of like, it’s taking a long time.” Others mentioned that automating coding could limit human error and help them optimize for inter-rater reliability. P13 coordinated a large project with multiple coders, and they described that “having a more thoughtful upfront conversation about that, and then using a tool to execute on that [code scheme], I think, would be way more accurate and better science than what we ended up doing.” Of those interested in using LLMs for coding, researchers had conflicting perspectives about whether LLMs would best be used for an initial, exploratory round of coding, or whether they should be used to apply manually developed codes. Some specifically identified that LLM-aided coding could help with the first round of coding, or when initially getting a grip on the data. P13 thought LLMs could “save you from that first tedious round of coding and then you could maybe be looking at the more nuanced codes faster over time.” Others described the desire to use LLMs as a way to “gut check” their coding. For example, P18 explained that they had considered if “me and another human are coding the data, and then ChatGPT is also coding the data, and we do an intercoder reliability check with all of us.” Nonetheless, some participants (P2, P4, P8, P16) struggled to see the automation of coding as aligned with perception of qualitative work, stating “there’s a part of qualitative research where it kind of feels like I’m in this for the manual work and that’s why I’m doing it” (P2).
@@ 4.1.3. LLMs in Other Research Tasks
Beyond data collection and analysis, participants also reported leveraging the interactivity of readily available LLMs, particularly ChatGPT. Participants felt that ChatGPT could help them ideate better and be more creative researchers; “I think LLMs are really great for bouncing ideas off of and thinking creatively,” P18 said, while P16 said they “will use it as like a basis or a starting point.” Participants also explained that they used ChatGPT to generate explanations for complicated concepts or coding implementations, like P7, who explained that they ask ChatGPT to help explain specific theories, stating “if I didn’t understand what kind of methods they used in a particular paper or […] I’m not able to make some connections or linkages, I just pose a question to ChatGPT.” In these instances, the social semblance of LLMs takes center stage, and our participants would sometimes refer to ChatGPT as more of a collaborator or peer than a tool. For example, P18 referred to ChatGPT as a “great brainstorming buddy,” P20 called it a “thought partner”, P9 defined it as “a rational person, a smart layman, if you will,” and P12 said that working with ChatGPT was “almost like working with another person.” For P11, this connection went further, as they found that interacting with ChatGPT while doing research helped stave off loneliness:
“As a researcher, we are always heads down and maybe we do not have much interaction with peer researchers who are human, o I feel like I really value this kind of interaction, somebody I can talk to. I feel like this research process is not boring, it’s engaging.”
@@
4.1.4. Viewing LLMs and Generative AI through the Lens of ChatGPT
Overall, perceptions of LLMs were strongly colored by ChatGPT’s performance as the most popular interface through which to explore LLMs. For example, this conflation impacted people’s general perception of LLMs’ suitability for complex tasks. Some participants tried to use ChatGPT for data annotation, observed mediocre results (“vague” (P6), “too generic” (P5), “very general” (P13), or “uninteresting” (P9)), and thus reached the conclusion that LLMs could not help their use case. However, some participants looked beyond out-of-the-box ChatGPT. For example, P17 set up a customized labeling pipeline that required some engineering, and P11 used a personal version of ChatGPT to improve results. By the same token, participants’ ideas about what might now be possible for the next generation of data analysis tools were likely expanded by ChatGPT and imagined natural language interfaces for tasks like data transformation and retrieval without technical knowledge. Generally speaking, participants felt that LLMs were more advanced in areas where ChatGPT performs better, thanks to their interactions directly through its web interface – for example, more participants felt that LLMs were helpful “brainstorming buddies” (P20), than felt they were helpful annotators, which is a use case for which ChatGPT in the web interface was not designed. Similarly, people’s perceptions of the privacy implications of sharing participant data with LLMs more broadly were influenced by their knowledge of ChatGPT and OpenAI. P16 demonstrated this tension nicely when they advised, “Don’t put anything your participants have said into LLMs. I don’t know much about it, but I’m guessing it collects the data.” This is an example of how multiple participants reported that they believed it unethical to share participant data with a third-party tool like ChatGPT, but did not discuss the potential alternative of open-source models to address some concerns. This may be due to a combined lack of knowledge on the subject or the relative scarcity of practically implemented alternatives for users, particularly those without technical backgrounds.
@@ 4.2.1. Uneven adoption
Researchers reported logistical hurdles to adopting AI tools, which seemed related to their comfort with technology and inequality around funding. For example, P13 expressed that the prevalent feeling among their team when faced with a new tool is: “We’re going to have to learn this, it’s gonna take forever, and we’re never going to get around to doing the actual project.” P2 and P5 recounted that cost influenced their choice of models, P2 stating “since the Claude model is quite expensive, I didn’t use it, and OpenAI models are much cheaper, so I could test it out.” Researchers also may not be certain how to find tools that meet their needs; P20 described how “you don’t know what’s out there, sometimes you don’t even know what something’s called.” Overall, logistical issues led researchers to be more reliant on accessible interfaces like ChatGPT, even when these are not optimized for their use case, an issue exacerbated for non-technical researchers.

### v2

@@ 4. Findings
In these findings, we outline how the researchers we interviewed adopted LLMs into qualitative research, identifying their main risks and concerns, then concluding by reflecting on the fundamental philosophical tensions raised.
@@ 4.1. Current and Potential Uses of Qualitative Researchers Adopting LLMs
We sought to understand how our participants had or had not experimented with using LLMs in their qualitative research process. Of the researchers we interviewed, fourteen reported actively exploring the use of LLMs in data collection or analysis. The most common use cases mentioned, highlighted in Figure 1, were using LLMs to help generate recruitment materials, attempts to use LLMs to speed up qualitative coding, and using LLMs for ideation and feedback.
ChatGPT was the most frequently mentioned LLM tool by participants. When we opened the conversation by asking participants about their experiences with “generative AI tools or large language models,” most would immediately talk about ChatGPT. Some researchers specified the model they used; P11 opened with “I use LLMs, specifically ChatGPT.” Others spoke more hesitantly, such as P13: “I’ve started using ChatGPT, I’m trying to be more familiar with it because it is the future, and you don’t want to be a dinosaur.” Overall, perceptions of LLMs writ large were strongly influenced by experiences with ChatGPT as the most popular interface through which to explore LLMs. Participants felt that LLMs were more advanced in areas where ChatGPT performed better, thanks to their interactions directly through the web interface. For example, more participants felt that LLMs were helpful “brainstorming buddies” (P20), than felt they were helpful annotators, a use case for which ChatGPT’s web interface was not designed.
Outside of ChatGPT, participants shared that they used a suite of available tools for qualitative research, such as NVivo, Atlas.ti, MAXQDA or Dedoose, and AI sometimes underpins these tools’ features. Participants were sometimes unsure whether these tools leveraged LLM-based functionality. For example, participants (P4, P8, P9, P13) wondered: “does Grammarly count?” Others mentioned explicitly that they engaged in AI features offered by qualitative coding platforms. For example, P7 reported using the in-context code suggestions that Atlas.ti provides. Participants did not necessarily feel agency over the AI features embedded in these tools, as P1 explained: “Atlas sends me a thematic analysis that essentially aggregates stuff, I don’t really like it,” and P4 summarized AI features in these tools as something they see “not by choice.” Overall, participants often relied on software platforms to help with qualitative research, though not all appreciated the apparent increase in “smart” AI-based features.
In contrast, some researchers shared they purposefully avoid advanced technology in their qualitative research process. While all participants used their computers for analysis, they also described paring back their approach to remain close to the data, particularly to data from participant interviews. P6 explained that while they were “not quite as old school as the people who would cut out their coded texts and move it around, I basically do that in Word documents.” P2 echoed this sentiment: “I didn’t like the restrictions and things the platform was putting on me, and I got really frustrated; I’m just familiar with Excel.” These attitudes highlight how, to some researchers, complex software interferes with their process or can be hard to learn, and simpler solutions may meet their needs.
@@ 4.1.1. LLMs during Data Collection
Seven participants reported using LLMs to help with creating or editing recruitment or interview materials. Often, this assistance took the form of researchers using a service like ChatGPT to edit or tweak their writing. Even those who were reluctant to use LLMs for research conceded that they could be helpful for writing, such as P10 who said “the only real reason that I use ChatGPT is to have it reword or rephrase,” and P4 explained that “it will give me a really good sentence or switch a word.” This capability was particularly helpful for researchers whose native language is not English; for example, P8 said they “feel more confident once [they] have passed [their] writing to the AI tools.” LLMs also assisted participants in tailoring recruitment materials to their subject population. P9 (a non-American, non-native English speaker) told us they used LLMs to make messaging more appropriate for American research participants. They shared, “one thing I noticed that LLMs can do beautifully is make things American,” and therefore “an LLM’s general answer or general email to an American future participant will probably be better than my take on it.” P5 reported that LLMs helped them deploy recruitment materials in multiple languages, explaining that they “used it for the translation of interviews and consent forms and other required documents,” before having a native speaker collaborator double check the output. LLMs were also used by participants to help them create interview and recruitment materials from scratch. For example, P8 recalled “I was working with a deadline, and I was feeling stuck. I was not sure what kinds of questions to ask, but I did have a set of broad topics […], so I put those into ChatGPT and I was like, ‘Can you help me come up with this type of interview question?”’ While quite a few of our participants had experimented with using these tools to help generate research materials, not all found these tools helpful: “the performance was really great,” P5 said, while P8 thought “I feel like I could have done better,” and P10 elaborated that its outputs were “really broad, versus the empirical work I’ve done is very specific and granular.”
@@
In terms of promising potential uses, researchers seemed to understand why one might use LLMs to help create recruitment materials. Some experienced researchers explained that they had past recruitment templates already, but for those who do not, P10 said, “I can imagine that maybe using something like this would be helpful for beginner researchers.” P20 expressed a similar attitude: “I would consider it in the future, if I’m drafting a new consent letter or things like that, but we had all of that from previous ones.” Nonetheless, some researchers pointed to specific reasons why they may not consider LLMs appropriate for these use cases, such as when materials are being produced for at-risk or non-Western populations, issues which will be further discussed in the following sections.
@@ 4.1.2. LLMs during Data Analysis
Another avenue that participants had explored, though often less successfully, was using an LLM to process and make sense of their qualitative data. Many researchers we interviewed did not attempt to use LLMs in data analysis due to concerns about privacy and performance, issues we explore in the next section. Among those interested in LLM-facilitated analysis of qualitative data, we saw various functions being performed with LLMs: transforming or retrieving data, data exploration, generating initial codes or themes, and applying qualitative codes.
@@
Researchers wanted to use LLMs as assistants for retrieving, processing, transforming, or reformatting their data using natural language. P9 reported that they leverage LLMs to perform fuzzy searches on qualitative data: “With retrieval via an LLM, I can say something like, ‘Where in this document is an example of such and such?’ and just talk more conceptually or more abstractly, and the tool will find it.” Participants also saw LLMs as being potentially useful to reformat data. For example, P4 imagined that “it would be nice if there was like a large language model agent to grab all the qualitative codes [and put them] onto a Google Jamboard because I have to rely on copy pasting.”
@@
On the analysis front, some participants explained that they valued being able to use LLM-aided software to flexibly explore participant data. In this data exploration use case, participants reported using chat-based models like ChatGPT to conduct initial sensemaking with and about their data. In P15’s words, the ability to easily analyze qualitative data with an LLM was a “game changer,” since “there’s no friction between the person trying to make sense of a lot of data and the information.” One participant said they once used ChatGPT to help them make sense of the theory behind their qualitative data, explaining that they had interviewed people about their experiences with grief. They were “getting all this lovely, well-gelled data where people were repeatedly saying the same things, everything was perfectly categorized” but the researcher was “having a hard time naming those categories in a way that was as inclusive as it could be.” In this example, the participant asked ChatGPT to come up with a list of potential themes, based not on their participant data but on prior theories, and it gave them “this list of things, and I was like, hmm, this is really neatly packaged […] it’s giving me labels for the things I’m seeing that are also very intuitive.” These themes then served as a jumping off point for deeper engagement with the data, and guided the researchers’ thinking for constructing their final qualitative codebook.
@@
Experiences were more mixed for participants who directly tried to use LLMs to generate or assign qualitative codes to data. In these use cases, participants were seeking to replace or augment the process of generating and/or applying a codebook to label qualitative data. The most common experience was perhaps that of P5, who “tried to use ChatGPT to get the very general theme and the results were not that great, so I decided to code it manually, but I tried it.” P9 engaged in a process they refer to as “semi-coding,” where they provided ChatGPT a chunk of text they already had intuitions about, and wanted to see if the LLM agreed with their perspective. Ultimately, however, they felt that the codes were “uninteresting in a way, or at least they were first level codes.” This experience was also shared by P11, who explained that when they “asked it to assign some initial codes [on their dataset], oftentimes the response will be very high level and abstract, so it’s not trustworthy at all.” P17 disagreed, and had set up specific coding infrastructure that allowed them to identify whether an individual piece of social media data contains a rare occurrence of interest to their research. P17 was seeking to identify and code specific types of gender-based online harassment in Reddit comments and derived the most advanced annotation pipeline we saw:
@@
In their workflow, P17 derived codes based on initial data, and then constructed an LLM pipeline to annotate each post, using the LLM to take a first pass at organizing the data. P17 expressed that this process does not preclude deep qualitative work, since they still do a manual layer of qualitative analysis.
Despite mixed current experiences using ChatGPT for data analysis, many participants felt curious or excited by the prospect of using LLMs to automate qualitative coding. The biggest factor that motivated the use of LLMs for annotation was the tedium and time investment required by manual coding. P10 said “I’m currently in the process of coding, and sometimes I’m kind of like, it’s taking a long time.” Others mentioned that automating coding could limit human error and help them optimize for inter-rater reliability. P13 coordinated a large project with multiple coders, and they described that “having a more thoughtful upfront conversation about that, and then using a tool to execute on that [code scheme], I think, would be way more accurate and better science than what we ended up doing.” Of those interested in using LLMs for coding, researchers had conflicting perspectives about whether LLMs would best be used for an initial, exploratory round of coding, or whether they should be used to apply manually developed codes. Some specifically identified that LLM-aided coding could help with the first round of coding, or when initially getting a grip on the data. P13 thought LLMs could “save you from that first tedious round of coding and then you could maybe be looking at the more nuanced codes faster over time.” Others described the desire to use LLMs as a way to “gut check” their coding. For example, P18 explained that they had considered if “me and another human are coding the data, and then ChatGPT is also coding the data, and we do an intercoder reliability check with all of us.” Nonetheless, some participants (P2, P4, P8, P16) struggled to see the automation of coding as aligned with their perception of qualitative work, stating “there’s a part of qualitative research where it kind of feels like I’m in this for the manual work and that’s why I’m doing it” (P2). Some perceived LLM involvement in some research tasks as fundamentally in tension with qualitative research values, which we will further revisit in a later section.
@@ 4.1.3. LLMs in Other Research Tasks
Beyond data collection and analysis, participants also reported leveraging the interactivity of readily available LLMs, particularly ChatGPT. Participants felt that ChatGPT could help them ideate better and be more creative researchers: “I think LLMs are really great for bouncing ideas off of and thinking creatively,” P18 said, while P16 said they use LLM outputs “a basis or a starting point.” Participants also explained that they used ChatGPT to generate explanations for complicated concepts or coding implementations, like P7, who explained that they ask ChatGPT to help explain specific theories, stating “if I didn’t understand what kind of methods they used in a particular paper or […] I’m not able to make some connections or linkages, I just pose a question to ChatGPT.” In these instances, the social semblance of LLMs takes center stage, and our participants would sometimes refer to ChatGPT as more of a collaborator or peer than a tool. For example, P18 referred to ChatGPT as a “great brainstorming buddy,” P20 called it a “thought partner”, P9 defined it as “a rational person, a smart layman, if you will,” and P12 said that working with ChatGPT was “almost like working with another person.” For P11, this connection went further, as they found that interacting with ChatGPT while doing research helped them stay engaged with their research:
“As a researcher, we are always heads down and maybe we do not have much interaction with peer researchers who are human, I feel like I really value this kind of interaction, somebody I can talk to. I feel like this research process is not boring, it’s engaging.”
@@ 4.2.1. Uneven adoption
Researchers reported logistical hurdles to adopting AI tools, which related to their comfort with technology and inequality around funding. For example, participants (P1, P3, P10, P13, P18) expressed that adopting new technologies takes time and effort: “We’re going to have to learn this, it’s gonna take forever, and we’re never going to get around to doing the actual project” (P13). These efforts could be even higher for researchers not in computing fields. Researchers also may not be certain how to find tools that meet their needs; P20 described how “you don’t know what’s out there, sometimes you don’t even know what something’s called.” P2 and P5 recounted that cost influenced their choice of models, P2 stating “since the Claude model is quite expensive, I didn’t use it, and OpenAI models are much cheaper, so I could test it out.” Overall, logistical issues led researchers to be more reliant on readily available interfaces like ChatGPT, even when these are not optimized for their use case, an issue exacerbated for non-technical researchers.